OpenStudy (jasmineflvs):

would x-21=2x simplified be 3x=21?

OpenStudy (jasmineflvs):

i think it is

OpenStudy (uri):

x-21=2x x-2x=21 -x=21 x=-21 :o

OpenStudy (ivettef365):

uri is right

OpenStudy (jasmineflvs):

okay

OpenStudy (uri):

When i'm not? c: jkkk

OpenStudy (ivettef365):

:)